The perimeter of a rectangle is 108m. The length is 3m more than twice the width. find the dimensions. What is the length in meters and what is the width in meters?
Answer by checkley77(12844)

L=2W+3
2L+2W=108
2(2W+3)+2W=108
4W+6+2W=108
6W=108-6
6W=102W=102/6
W=17 M IS THE WIDTH.
L=2*17+3=34+3=37 M IS THE LENGTH.
PROOF:
2*37=2*17=108
74+34=108
108=108